In laser eye surgery, the laser emits a 1.49-nspulse focused on a spot that is 33.0 μm in diameter.

Part A

If the energy contained in the pulse is 2.85 mJ, what is the power per square meter (the irradiance) associated with this beam?

Express your answer to three significant figures.

I =   W/m2  

Part B

Suppose a molecule with a diameter of 0.500 nm is irradiated by the laser beam. How much energy does the molecule receive in one pulse from the laser? (The energy obtained in this part is more than enough to dissociate a molecule.)

Express your answer to three significant figures.

Emol =   MeV
Answer #1

a )

the power per square meter (the irradiance) associated with this beam is " P "

i = E / A = E / \pi ( d/2 )2

i = 2.85 X 10-3 X / 3.14 X ( 33 X 10-6 / 2 )2  

i = 3.333 X 106 W/m2

b )

Emol = i A

= 3.333 X 106 X 3.14 X ( 0.5 X 10-9 / 2 )2  

Emol = 6.541 X 10-13 J

Emol = 6.541 X 10-13 / 1.6 X 10-13 MeV

Emol = 4.088 MeV